# Procedure Learning via Regularized Gromov-Wasserstein Optimal Transport

### v1

@@ 3.1.2 Degenerate Solution
Although the FGWOT formulation has demonstrated great performance on several video alignment benchmarks, including Pouring, Penn Action, and IKEA ASM, in [3], optimizing exclusively for the FGWOT objective may potentially lead to trivial solutions. More specifically, since the cost matrix $\mathbf{C}$ minimizes visual differences between corresponding frames in $X$ and $Y$, and both the temporal prior $\mathbf{R}$ and structural priors $\mathbf{C}^{x}$ and $\mathbf{C}^{y}$ minimizes temporal differences between corresponding frames in $X$ and $Y$ (e.g., temporally nearby frames in $X$ should be paired with temporally adjacent frames in $Y$ and vice versa), there is no mechanism to prevent the optimization from collapsing to degenerate solutions, where all frames are mapped to a small cluster in the embedding space, and hence an entire video is assigned to a single key step. We consistently observe this issue on ProceL and illustrate some qualitative results in Fig. 3. From the results, every video is associated with only one key step, whereas the ground truth contains many. We note that due to different dataset characteristics, this problem seldom appears on EgoproceL and CrossTask.
Figure 3: Degenerate solutions by [3] across four different subtasks of ProceL [23]. Ground truth and results by [3] are shown at top and bottom rows respectively.
@@ 3.1.3 Contrastive Regularization
To avoid trivial solutions, we incorporate Contrastive Inverse Difference Moment (C-IDM) [31] as a regularization, which is applied separately on frame embeddings $\mathbf{X}$ and $\mathbf{Y}$, yielding our regularized Gromov-Wasserstein optimal transport framework (RGWOT), as illustrated in Fig. 2. Since $\mathcal{L}_{reg}(\mathbf{X})$ and $\mathcal{L}_{reg}(\mathbf{Y})$ are similar, below we describe $\mathcal{L}_{reg}(\mathbf{X})$ only, which is written as:
@@ 3.2 Clustering and Key-Step Ordering
Given frame embeddings learned via the above RGWOT approach, our next steps include localizing key steps and determining their order for procedure learning. We model key step localization as multi-label graph-cut segmentation [29], following [4]. We first build a graph which includes $K$ terminal nodes representing $K$ key steps and non-terminal nodes representing frame embeddings, as well as t-links connecting non-terminal nodes to terminal nodes and n-links connecting two non-terminal nodes. We employ $\alpha$-Expansion [8] to obtain the optimal cut, yielding frame assignments to $K$ key steps. Next, for each video, the order of key steps are determined by sorting the average normalized time of their assigned frames, following [4]. Finally, we rank the key-step orders based on their frequency appearing in the videos and output the top ranked key-step order.
@@ 4 Experiments
Datasets. We test our RGWOT approach on three datasets: one first-person dataset, EgoProceL [4], and two third-person datasets, ProceL [23] and CrossTask [74]. EgoProceL contains 62 hours of egocentric videos capturing 16 diverse tasks. ProceL contains 720 videos of 12 tasks with visually similar key-steps that span a total of 47 hours. CrossTask consists of 2763 videos totaling 213 hours for 18 tasks. Following previous works [4, 5, 11], the same training and validation splits were used for all methods.
Competing Methods. We compare our RGWOT method with previous self-supervised procedure learning methods, as well as with random and uniform distributions. Competing methods include CNC [4], GPL [5] and OPEL [11]. OPEL, in particular, uses a Kantorovich optimal transport formulation, making it our closest competitor.
Evaluation Metrics. We evaluate our RGWOT approach for procedure learning on the validation set of all datasets in accordance with prior works [4, 5, 11]. For the egocentric dataset (EgoProceL [4]), we report F1-score and Intersection over Union (IoU). Furthermore, to evaluate our RGWOT approach on third-person datasets (ProceL [23] and CrossTask [74]), we report precision, recall, and F1-score. Framewise scores are computed separately for each keystep, and the final score is reported as the average across all keysteps. We apply the Hungarian matching algorithm [40] to align the predicted labels with the ground truth, similarly to prior works  [23, 22, 59, 4, 5, 11].
Implementation Details. We adopt ResNet-50  [32] as the backbone network for feature extraction. We trained the encoder network using a pair of videos, randomly sampling frames from each during training as done by [21]. The network is optimized with the proposed RGWOT loss. Features are taken from the Conv4c layer, and a sequence of $c$ context frame features is stacked along the temporal axis. To encode temporal information, the stacked features are fed into two 3D convolutional layers, followed by a 3D global max pooling layer. The resulting representation is then passed through two fully connected layers and finally projected into a 128-dimensional embedding space using a linear layer. We follow the task-specific protocols from [22, 4, 11] in all experiments.
@@ 4.1 State-of-the-Art Comparison Results
Comparisons on Egocentric Dataset. Tab. 1 provides comparative evaluation of our RGWOT approach against previous methods on the large-scale egocentric (first-person) benchmark of EgoProceL [4]. In particular, it is a recent dataset tailored for egocentric procedure learning, serving as a strong benchmark for evaluating new methods on egocentric videos. Remarkably, RGWOT consistently surpasses existing approaches across all sub-datasets of EgoProceL [4]. Notably, RGWOT achieves an average improvement of 15.1% in F1 score and 15.3% in IoU compared to previous methods. These results demonstrate the superiority of our RGWOT formulation with a structural prior over OPEL [11]’s traditional Kantorovich optimal transport formulation with an optimality prior.
Table 1: Comparisons on egocentric dataset (i.e., EgoProceL [4]). Best results are in bold, while second best are underlined.
EgoProceL [4]
CMU-MMAC [15] EGTEA-GAZE+ [46] MECCANO [56] EPIC-Tents [35] PC Assembly PC Disassembly
@@
CnC [4] 22.7 11.1 21.7 9.5 18.1 7.8 17.2 8.3 25.1 12.8 27.0 14.8
GPL-2D [5] 21.8 11.7 23.6 14.3 18.0 8.4 17.4 8.5 24.0 12.6 27.4 15.9
UG-I3D [5] 28.4 15.6 25.3 14.7 18.3 8.0 16.8 8.2 22.0 11.7 24.2 13.8
GPL-w BG [5] 30.2 16.7 23.6 14.9 20.6 9.8 18.3 8.5 27.6 14.4 26.9 15.0
GPL-w/o BG [5] 31.7 17.9 27.1 16.0 20.7 10.0 19.8 9.1 27.5 15.2 26.7 15.2
OPEL [11] 36.5 18.8 29.5 13.2 39.2 20.2 20.7 10.6 33.7 17.9 32.2 16.9
@@
Comparisons on Third-Person Datasets. In Tab. 2, we evaluate the performance of our RGWOT approach against previous self-supervised procedure learning models on two third-person datasets, i.e., ProceL [23] and CrossTask [74]. It is evident that RGWOT achieves the best overall results across both datasets, outperforming all existing models. Notably, RGWOT demonstrates a significant improvement over OPEL [11], the previous best performing approach. Our approach surpasses OPEL [11] by 9.4% on ProceL [23] and by 5.4% on CrossTask[74] on the F1 scores.
Table 2: Comparisons on third-person datasets (i.e., ProceL [23] and CrossTask [74]). Best results are in bold, while second best are underlined.
ProceL [23] CrossTask [74]
@@
Alayrac et al. [2] 12.3 3.7 5.5 6.8 3.4 4.5
Kukleva et al. [41] 11.7 30.2 16.4 9.8 35.9 15.3
Elhamifar et al. [22] 9.5 26.7 14.0 10.1 41.6 16.3
Fried et al. [26] 28.8
Shen et al. [59] 16.5 31.8 21.1 15.2 35.5 21.0
CnC [4] 20.7 22.6 21.6 22.8 22.5 22.6
GPL-2D [61] 21.7 23.8 22.7 24.1 23.6 23.8
UG-I3D [61] 21.3 23.0 22.1 23.4 23.0 23.2
GPL [61] 22.4 24.5 23.4 24.9 24.1 24.5
STEPs [58] 23.5 26.7 24.9 26.2 25.8 25.9
OPEL [11] 33.6 36.3 34.9 35.6 34.8 35.1
@@
Comparisons with Multimodal Method. Tab. 3 highlights the performance of our RGWOT approach compared to STEPs [58], a multimodal approach to unsupervised procedure learning. STEPs [58] utilizes depth and gaze data along with RGB while our RGWOT approach uses only RGB data. Our RGWOT approach outperforms STEPS [58] on most datasets. RGWOT only has lower F1 score on EPIC-Tents [35], while still outperforming it on IoU. It should also be noted that our approach outperforms models that utilize narrations along with video data, namely Alayrac et al. [2] and Shen et al. [59], as observed in Tab. 2.
Table 3: Comparisons against multimodal method on EgoProceL [4]. Best results are in bold, while second best are underlined.
CMU-MMAC [15] EGTEA-GAZE+ [46] MECCANO [56] EPIC-Tents [35] ProceL [23] CrossTask [74]
@@
Table 4: Comparisons against action segmentation methods on ProceL [23] and CrossTask [74]. Best results are in bold, while second best are underlined.
Action Segmentation Method ProceL [23] CrossTask [74]
@@
JointSeqFL [23] - - 29.8 - - -
Elhamifar et al. [22] 9.5 26.7 14.0 10.1 41.6 16.3
Fried et al. [26] - - - - 28.8 -
Shen et al. [59] 16.5 31.8 21.1 15.2 35.5 21.0
Dvornik et al. [19] - - - - - 25.3
StepFormer [20] 18.3 28.1 21.9 22.1 42.0 28.3
OPEL [11] 33.6 36.3 34.9 35.6 34.8 35.1
@@
Tab. 4 presents a comparison of our RGWOT approach against several state-of-the-art unsupervised action segmentation methods, as well as OPEL [11], on the ProceL [23] and CrossTask [74] datasets. On the ProceL [23] dataset, RGWOT achieves a substantial improvement with the highest precision, recall, and F1 score, significantly surpassing previous methods. On the CrossTask [74] dataset, although StepFormer [20] achieves the highest recall, its F1 score remains relatively low at 28.3%, indicating a poor balance between precision and recall. In contrast, RGWOT achieves the highest F1 score of 40.4% with a well-balanced precision (40.4%) and recall (40.7%), which is critical since F1 provides a more reliable measure of overall segmentation quality. These results demonstrate that RGWOT is robust and effective, consistently outperforming other approaches across both datasets.
@@
Figure 4: Qualitative results on ProceL [23]. Colored segments represent predicted actions with a particular color denoting the same action across all models.
Fig. 4 illustrates example results of unsupervised procedure learning on two ProceL [23] videos. The CnC [4] model tends to over-segment, and has generally worse alignment for the clusters it predicts correctly. OPEL [11] fares better but it also over-segments and misaligns actions. In contrast, our RGWOT model aligns tasks significantly closer to the ground truth and avoids over-segmentation. This visualizes the superior procedure learning capabilities of RGWOT over previous works.
@@ 4.2 Ablation Study Results
Effects of Model Components. Tab. 5 presents an ablation study where we systematically remove one model component at a time to assess its contribution to the overall model performance. The baseline configuration includes all four components: contrastive regularization, temporal prior, structural prior, and virtual frame, achieving the highest results. Removing the virtual frame component leads to a substantial performance drop, indicating its critical role. The absence of the temporal and structural priors also results in moderate declines, underscoring their importance. Excluding contrastive regularization causes a smaller degradation, suggesting it is less important on EgoProceL [4]. Nevertheless, contrastive regularization is critical in preventing degenerate solutions on ProceL [23], as shown in Fig. 3. Overall, incorporating all components yields the most robust performance, thereby justifying their inclusion in the complete architecture.
Table 5: Effects of model components on performance across MECCANO [56] and PC Disassembly. Best results are in bold, while second best are underlined.
Contrastive Regularization Temporal Prior Structural Prior Virtual Frame MECCANO [56] PC Disassembly
@@
CMU-MMAC [15] EGTEA-GAZE+ [46] MECCANO [56] EPIC-Tents [35]
@@
Effects of Training Data Quantity. Fig. 5 shows how model performance improves as more training videos are used on the MECCANO [56] dataset. Our method, RGWOT, consistently outperforms previous methods across all training set sizes. Even with just two training videos, RGWOT achieves a strong F1 score, surpassing the best scores of the other methods. As more data is added, its performance increases steadily, reaching 59.5% with 17 videos. In contrast, previous methods exhibit slower gains and lower overall performance. These results highlight RGWOT’s superior data efficiency and scalability.
Figure 5: Effects of training data quantity on F1 score across different methods on MECCANO [56].
Supplementary Material. Due to space constraints, please refer to our supplementary material for our hyperparameter settings and quantitative results on all subtasks of EgoProceL [4], ProceL [23], and CrossTask [74].

### v2

@@ 3.1.2 Degenerate Solution
Although the FGWOT formulation has demonstrated great performance on several video alignment benchmarks, including Pouring, Penn Action, and IKEA ASM, in [ali2025joint], optimizing exclusively for the FGWOT objective may potentially lead to trivial solutions. More specifically, since the cost matrix $\mathbf{C}$ minimizes visual differences between corresponding frames in $X$ and $Y$, and both the temporal prior $\mathbf{R}$ and structural priors $\mathbf{C}^{x}$ and $\mathbf{C}^{y}$ minimizes temporal differences between corresponding frames in $X$ and $Y$ (e.g., temporally nearby frames in $X$ should be paired with temporally adjacent frames in $Y$ and vice versa), there is no mechanism to prevent the optimization from collapsing to degenerate solutions, where all frames are mapped to a small cluster in the embedding space, and hence an entire video is assigned to a single key step. We consistently observe this issue on ProceL and illustrate some qualitative results in Fig. 3. From the results, every video is associated with only one key step, whereas the ground truth contains many. We note that due to different dataset characteristics, this problem seldom appears on EgoproceL and CrossTask.
Figure 3: Degenerate solutions by [ali2025joint] across four different subtasks of ProceL [elhamifar2019unsupervised]. Ground truth and results by [ali2025joint] are shown at top and bottom rows respectively.
@@ 3.1.3 Contrastive Regularization
To avoid trivial solutions, we incorporate Contrastive Inverse Difference Moment (C-IDM) [haresh2021learning] as a regularization, which is applied separately on frame embeddings $\mathbf{X}$ and $\mathbf{Y}$, yielding our regularized Gromov-Wasserstein optimal transport framework (RGWOT), as illustrated in Fig. 2. Since $\mathcal{L}_{reg}(\mathbf{X})$ and $\mathcal{L}_{reg}(\mathbf{Y})$ are similar, below we describe $\mathcal{L}_{reg}(\mathbf{X})$ only, which is written as:
@@
Finally, our training objective is written as:
$\mathcal{L}=\mathcal{L}_{align}(\mathbf{X},\mathbf{Y})+\beta\left(\mathcal{L}_{reg}(\mathbf{X})+\mathcal{L}_{reg}(\mathbf{Y})\right),$ (7)
where $\beta$ is the weight for the contrastive regularization.
@@ 3.2 Clustering and Key-Step Ordering
Given frame embeddings learned via the above RGWOT approach, our next steps include localizing key steps and determining their order for procedure learning. We model key step localization as multi-label graph-cut segmentation [greig1989exact], following [bansal2022my]. We first build a graph which includes $K$ terminal nodes representing $K$ key steps and non-terminal nodes representing frame embeddings, as well as t-links connecting non-terminal nodes to terminal nodes and n-links connecting two non-terminal nodes. We employ $\alpha$-Expansion [boykov2002fast] to obtain the optimal cut, yielding frame assignments to $K$ key steps. Next, for each video, the order of key steps are determined by sorting the average normalized time of their assigned frames, following [bansal2022my]. Finally, we rank the key-step orders based on their frequency appearing in the videos and output the top ranked key-step order.
@@ 4 Experiments
Datasets. We test our RGWOT approach on three datasets: one first-person dataset, EgoProceL [bansal2022my], and two third-person datasets, ProceL [elhamifar2019unsupervised] and CrossTask [zhukov2019cross]. EgoProceL contains 62 hours of egocentric videos capturing 16 diverse tasks. ProceL contains 720 videos of 12 tasks with visually similar key-steps that span a total of 47 hours. CrossTask consists of 2763 videos totaling 213 hours for 18 tasks. Following previous works [bansal2022my, bansal2024united, chowdhuryopel], the same training and validation splits were used for all methods.
Competing Methods. We compare our RGWOT method with previous self-supervised procedure learning methods, as well as with random and uniform distributions. Competing methods include CNC [bansal2022my], GPL [bansal2024united] and OPEL [chowdhuryopel]. OPEL, in particular, uses a Kantorovich optimal transport formulation, making it our closest competitor.
Evaluation Metrics. We evaluate our RGWOT approach for procedure learning on the validation set of all datasets in accordance with prior works [bansal2022my, bansal2024united, chowdhuryopel]. For the egocentric dataset (EgoProceL [bansal2022my]), we report F1-score and Intersection over Union (IoU). Furthermore, to evaluate our RGWOT approach on third-person datasets (ProceL [elhamifar2019unsupervised] and CrossTask [zhukov2019cross]), we report precision, recall, and F1-score. Framewise scores are computed separately for each keystep, and the final score is reported as the average across all keysteps. We apply the Hungarian matching algorithm [kuhn1955hungarian] to align the predicted labels with the ground truth, similarly to prior works  [elhamifar2019unsupervised, elhamifar2020self, shen2021learning, bansal2022my, bansal2024united, chowdhuryopel].
Implementation Details. We adopt ResNet-50  [he2016deep] as the backbone network for feature extraction. We trained the encoder network using a pair of videos, randomly sampling frames from each during training as done by [dwibedi2019tcc]. The network is optimized with the proposed RGWOT loss. Features are taken from the Conv4c layer, and a sequence of $c$ context frame features is stacked along the temporal axis. To encode temporal information, the stacked features are fed into two 3D convolutional layers, followed by a 3D global max pooling layer. The resulting representation is then passed through two fully connected layers and finally projected into a 128-dimensional embedding space using a linear layer. We follow the task-specific protocols from [elhamifar2020self, bansal2022my, chowdhuryopel] in all experiments.
@@ 4.1 State-of-the-Art Comparison Results
Comparisons on Egocentric Dataset. Tab. 1 provides comparative evaluation of our RGWOT approach against previous methods on the large-scale egocentric (first-person) benchmark of EgoProceL [bansal2022my]. In particular, it is a recent dataset tailored for egocentric procedure learning, serving as a strong benchmark for evaluating new methods on egocentric videos. Remarkably, RGWOT consistently surpasses existing approaches across all sub-datasets of EgoProceL [bansal2022my]. Notably, RGWOT achieves an average improvement of 15.1% in F1 score and 15.3% in IoU compared to previous methods. These results demonstrate the superiority of our RGWOT formulation with a structural prior over OPEL [chowdhuryopel]’s traditional Kantorovich optimal transport formulation with an optimality prior.
Table 1: Comparisons on egocentric dataset (i.e., EgoProceL [bansal2022my]). Best results are in bold, while second best are underlined.
EgoProceL [bansal2022my]
CMU-MMAC [de2009guide] EGTEA-GAZE+ [li2018eye] MECCANO [ragusa2021meccano] EPIC-Tents [jang2019epic] PC Assembly PC Disassembly
@@
CnC [bansal2022my] 22.7 11.1 21.7 9.5 18.1 7.8 17.2 8.3 25.1 12.8 27.0 14.8
GPL-2D [bansal2024united] 21.8 11.7 23.6 14.3 18.0 8.4 17.4 8.5 24.0 12.6 27.4 15.9
UG-I3D [bansal2024united] 28.4 15.6 25.3 14.7 18.3 8.0 16.8 8.2 22.0 11.7 24.2 13.8
GPL-w BG [bansal2024united] 30.2 16.7 23.6 14.9 20.6 9.8 18.3 8.5 27.6 14.4 26.9 15.0
GPL-w/o BG [bansal2024united] 31.7 17.9 27.1 16.0 20.7 10.0 19.8 9.1 27.5 15.2 26.7 15.2
OPEL [chowdhuryopel] 36.5 18.8 29.5 13.2 39.2 20.2 20.7 10.6 33.7 17.9 32.2 16.9
VAOT [ali2025joint] 50.3 34.9 35.1 20.4 57.3 41.2 33.5 20.4 38.1 23.6 44.4 28.9
@@
Comparisons on Third-Person Datasets. In Tab. 2, we evaluate the performance of our RGWOT approach against previous self-supervised procedure learning models on two third-person datasets, i.e., ProceL [elhamifar2019unsupervised] and CrossTask [zhukov2019cross]. It is evident that RGWOT achieves the best overall results across both datasets, outperforming all existing models. Notably, RGWOT demonstrates a significant improvement over OPEL [chowdhuryopel], the previous best performing approach. Our approach surpasses OPEL [chowdhuryopel] by 9.4% on ProceL [elhamifar2019unsupervised] and by 5.4% on CrossTask[zhukov2019cross] on the F1 scores.
Table 2: Comparisons on third-person datasets (i.e., ProceL [elhamifar2019unsupervised] and CrossTask [zhukov2019cross]). Best results are in bold, while second best are underlined. ‘X’ denotes degenerate results.
ProceL [elhamifar2019unsupervised] CrossTask [zhukov2019cross]
@@
Alayrac et al. [alayrac2016unsupervised] 12.3 3.7 5.5 6.8 3.4 4.5
Kukleva et al. [kukleva2019unsupervised] 11.7 30.2 16.4 9.8 35.9 15.3
Elhamifar et al. [elhamifar2020self] 9.5 26.7 14.0 10.1 41.6 16.3
Fried et al. [fried2020learning] 28.8
Shen et al. [shen2021learning] 16.5 31.8 21.1 15.2 35.5 21.0
CnC [bansal2022my] 20.7 22.6 21.6 22.8 22.5 22.6
GPL-2D [si2020adversarial] 21.7 23.8 22.7 24.1 23.6 23.8
UG-I3D [si2020adversarial] 21.3 23.0 22.1 23.4 23.0 23.2
GPL [si2020adversarial] 22.4 24.5 23.4 24.9 24.1 24.5
STEPs [shah2023steps] 23.5 26.7 24.9 26.2 25.8 25.9
OPEL [chowdhuryopel] 33.6 36.3 34.9 35.6 34.8 35.1
VAOT [ali2025joint] X X X 38.2 38.9 38.6
@@
Table 3: Comparisons against multimodal method on EgoProceL [bansal2022my]. Best results are in bold, while second best are underlined.
CMU-MMAC [de2009guide] EGTEA-GAZE+ [li2018eye] MECCANO [ragusa2021meccano] EPIC-Tents [jang2019epic] ProceL [elhamifar2019unsupervised] CrossTask [zhukov2019cross]
@@
Comparisons with Multimodal Method. Tab. 3 highlights the performance of our RGWOT approach compared to STEPs [shah2023steps], a multimodal approach to unsupervised procedure learning. STEPs [shah2023steps] utilizes depth and gaze data along with RGB while our RGWOT approach uses only RGB data. Our RGWOT approach outperforms STEPS [shah2023steps] on most datasets. RGWOT only has lower F1 score on EPIC-Tents [jang2019epic], while still outperforming it on IoU. It should also be noted that our approach outperforms models that utilize narrations along with video data, namely Alayrac et al. [alayrac2016unsupervised] and Shen et al. [shen2021learning], as observed in Tab. 2.
@@
Table 4: Comparisons against action segmentation methods on ProceL [elhamifar2019unsupervised] and CrossTask [zhukov2019cross]. Best results are in bold, while second best are underlined.
Action Segmentation Method ProceL [elhamifar2019unsupervised] CrossTask [zhukov2019cross]
@@
JointSeqFL [elhamifar2019unsupervised] - - 29.8 - - -
Elhamifar et al. [elhamifar2020self] 9.5 26.7 14.0 10.1 41.6 16.3
Fried et al. [fried2020learning] - - - - 28.8 -
Shen et al. [shen2021learning] 16.5 31.8 21.1 15.2 35.5 21.0
Dvornik et al. [dvornik2022flow] - - - - - 25.3
StepFormer [dvornik2023stepformer] 18.3 28.1 21.9 22.1 42.0 28.3
OPEL [chowdhuryopel] 33.6 36.3 34.9 35.6 34.8 35.1
@@
Tab. 4 presents a comparison of our RGWOT approach against several state-of-the-art unsupervised action segmentation methods, as well as OPEL [chowdhuryopel], on the ProceL [elhamifar2019unsupervised] and CrossTask [zhukov2019cross] datasets. On the ProceL [elhamifar2019unsupervised] dataset, RGWOT achieves a substantial improvement with the highest precision, recall, and F1 score, significantly surpassing previous methods. On the CrossTask [zhukov2019cross] dataset, although StepFormer [dvornik2023stepformer] achieves the highest recall, its F1 score remains relatively low at 28.3%, indicating a poor balance between precision and recall. In contrast, RGWOT achieves the highest F1 score of 40.4% with a well-balanced precision (40.4%) and recall (40.7%), which is critical since F1 provides a more reliable measure of overall segmentation quality. These results demonstrate that RGWOT is robust and effective, consistently outperforming other approaches across both datasets.
@@
Figure 4: Qualitative results on ProceL [elhamifar2019unsupervised]. Colored segments represent predicted actions with a particular color denoting the same action across all models.
Fig. 4 illustrates example results of unsupervised procedure learning on two ProceL [elhamifar2019unsupervised] videos. The CnC [bansal2022my] model tends to over-segment, and has generally worse alignment for the clusters it predicts correctly. OPEL [chowdhuryopel] fares better but it also over-segments and misaligns actions. In contrast, our RGWOT model aligns tasks significantly closer to the ground truth and avoids over-segmentation. This visualizes the superior procedure learning capabilities of RGWOT over previous works.
@@ 4.2 Ablation Study Results
Effects of Model Components. Tab. 5 presents an ablation study where we systematically remove one model component at a time to assess its contribution to the overall model performance. The baseline configuration includes all four components: contrastive regularization, temporal prior, structural prior, and virtual frame, achieving the highest results. Removing the virtual frame component leads to a substantial performance drop, indicating its critical role. The absence of the temporal and structural priors also results in moderate declines, underscoring their importance. Excluding contrastive regularization causes a smaller degradation, suggesting it is less important on EgoProceL [bansal2022my]. Nevertheless, contrastive regularization is critical in preventing degenerate solutions on ProceL [elhamifar2019unsupervised], as shown in Fig. 3. Overall, incorporating all components yields the most robust performance, thereby justifying their inclusion in the complete architecture.
Table 5: Effects of model components on performance across MECCANO [ragusa2021meccano] and PC Disassembly. Best results are in bold, while second best are underlined.
Contrastive Regularization Temporal Prior Structural Prior Virtual Frame MECCANO [ragusa2021meccano] PC Disassembly
@@
CMU-MMAC [de2009guide] EGTEA-GAZE+ [li2018eye] MECCANO [ragusa2021meccano] EPIC-Tents [jang2019epic]
@@
Effects of Training Data Quantity. Fig. 5 shows how model performance improves as more training videos are used on the MECCANO [ragusa2021meccano] dataset. Our method, RGWOT, consistently outperforms previous methods across all training set sizes. Even with just two training videos, RGWOT achieves a strong F1 score, surpassing the best scores of the other methods. As more data is added, its performance increases steadily, reaching 59.5% with 17 videos. In contrast, previous methods exhibit slower gains and lower overall performance. These results highlight RGWOT’s superior data efficiency and scalability.
Figure 5: Effects of training data quantity on F1 score across different methods on MECCANO [ragusa2021meccano].
Supplementary Material. Due to space constraints, please refer to our supplementary material for our hyperparameter settings, sensitivity analyses, run time comparisons, degenerate solution justification, and quantitative results on all subtasks of EgoProceL [bansal2022my], ProceL [elhamifar2019unsupervised], and CrossTask [zhukov2019cross].